OpenStudy (anonymous):

-10x=1/2(8x+14) <-- solve and check the algebraic expression , please show work

OpenStudy (anonymous):

1) Simplify it (see the attachment) 2) -10x = 4x + 7 so 14x = -7 x = -1/2

OpenStudy (anonymous):

i did this -10x=1/2(8x+14) -10x=4x+7 -7 -7 ------------ -17x=4x <--- idk if i did something wrong or whats the next step , and how would i plug it in so i can check the answer is correct ?

OpenStudy (anonymous):

when you subtract both side by -7, it's not -7X.... (not an x-term) so you can't do this: -10x -7x should be -10x -7